An research exploration company currently has two active projects, one in Africa and the other in Europe. The probability that the European project is successful is 0.3 more than that of the African project. There is a 28% chance that both projects will be successful and an 18% chance that neither project will be successful. What is the probability that the European project will be successful?
Answer #1

let probability of European and African project to be successful are P(E) and P(A)

from above P(E)-P(A) =0.3...........(1)

P(E n A) =0.28 ...........(2)

P(E U A) =P(E)+P(A)-P(E n A) =1-0.18

P(E)+P(A) -0.28 =0.82

P(E)+P(A) =1.10 .........(3)

solving equation (1) and (2)

P(E) =(0.3+1.1)/2 =0.7

probability that the European project will be successful =0.7